# Leveraging Quantum Superposition to Infer the Dynamic Behavior of a Spatial-Temporal Neural Network Signaling Model

### v2

@@ 7 Discussion
In this work, we introduced a novel class of problems related to the dynamics of large-scale networks relevant to neuroscience and machine learning. The problem we addressed asks whether a network can sustain inherent dynamic activity, cease functioning through quiescence, or saturate in an epileptic-like state. By carefully structuring the problem, we were able to take advantage of the unique properties of quantum computing, particularly the Deutsch–Jozsa and Grover algorithms, and provide an approach that solves this problem with significant computational efficiency compared to classical methods.
@@
The importance of this work lies in both its theoretical and practical implications. For neuroscience, it provides a potential novel tool to better model and understand the temporal and structural properties of neural networks. More broadly, it has potential applications in analyzing how brain dynamics relate to disorders associated with network disconnections, such as epilepsy, neurodevelopmental disorders such as Autism Spectrum Disorder, and neural cognitive and degenerative disorders. In the context of machine learning and artificial intelligence, it is relevant to constructing and optimizing artificial neural network models. For quantum computing, this work serves as a proof-of-concept for how existing quantum algorithms can be adapted and extended to solve new, non-trivial interdisciplinary problems, motivating the continued exploration and expansion of quantum algorithms and computing.
7.1 Modeling the Evolution of Network Dynamics Beyond $T_{o}$
The framework and approach we introduce in this paper necessitate time evolving the network dynamics until the observation time $T_{o}$. A natural criticism is that if the dynamics can be evolved classically to reach $T_{o}$, it may be computationally simpler to continue evolving the system classically beyond $T_{o}$ rather than leveraging a quantum approach. While this is a reasonable point, our intent here, in part, was to show that certain classes of problems related to network dynamics can be formally structured to give a quantum approach solution a computational advantage. In effect, it is a proof of concept example intended to motivate future work.
However, this consideration raises two broad areas in which quantum computing may contribute to the scientific study of neural networks. The first is large-scale simulations of neural dynamics across scales of spatial and temporal organization, bounded and informed by the known physiology (in the case of the brain) and known models (for artificial intelligence and machine learning). This is beyond the intent or scope of this paper. The second is carefully chosen and structured problems about neural dynamics that use quantum algorithms carefully and in a clever way, which we addressed in the current work. We propose they are necessarily complimentary rather than mutually exclusive or distinct.
@@ 7.1 Modeling the Evolution of Network Dynamics Beyond $T_{o}$
There may also be scenarios where, instead of a network model being simulated, a network evolves naturally according to its physical makeup, i.e., activity in the brain or internal activation patterns in artificial neural networks. In this scenario, only then are the dynamics measured or observed at some arbitrary time $T_{o}$ to probe its behavior. In this case, as we showed here, a quantum computational approach to a well-structured question about the dynamics could be the only way to answer the question. This could have very real-world implications. The dynamics of the network may be inaccessible or uninterpretable before or beyond $T_{o}$. For example, in biological neural networks, experimental limitations, such as noise, finite measurement windows, or ethical considerations, may constrain observations to specific time windows.
@@
7.2 Open Questions and Future Directions
While our work establishes a foundational framework, it also raises several important questions and opportunities for future research. First and foremost, while we showed a novel application of quantum computing to neural network dynamics and demonstrated significant computational speedup theoretically, implementing the approach we developed on current quantum hardware presents many practical challenges. Noise, decoherence, and the limited number of logical qubits in today’s quantum processors remain prohibitive limitations to modeling or computing properties of networks of any meaningful size. Future work could focus on testing our approach on new hardware and even adopt this as a benchmark problem.
We used the Deutsch–Jozsa and Grover algorithms in this work intentionally in part because we wanted to see if we could structure and solve a meaningful neural network problem by leveraging two foundational and extensively studied quantum algorithms. However, other quantum algorithms may offer complementary or even enhanced capabilities for similar types of problems. For example, algorithms designed for graph analysis or hybrid quantum-classical approaches could further expand the scope of network dynamics problems solvable by quantum computing.
In this paper and almost all of our prior published work on the neural dynamic model, we focused on fractional summation but deterministic network dynamics. However, in particular, neurons and neurobiological networks are inherently noisy and probabilistic. Extending the framework to account for stochastic dynamics is likely to be an important algorithmic strategy used by nature, but it raises new challenges about how to compute probabilistic states on quantum devices.
Lastly, in an even broader context, the methods developed here could inspire the design and engineering of quantum neural networks relevant to quantum machine learning that incorporate dynamic feedback mechanisms similar to biological neurons. Such networks could have applications in learning systems where temporal patterns and adaptive dynamics are critical.

### v4

@@ 8 Discussion
In this work, we introduced a particular class of problem related to the dynamics of large-scale networks relevant to neuroscience and machine learning. The problem we addressed asks whether a network can sustain inherent dynamic activity, cease functioning through quiescence, or saturate in an epileptic-like state. By carefully structuring the problem, we were able to take advantage of the unique properties of quantum computing, particularly the Deutsch–Jozsa and Grover algorithms, and provide an approach that solves this problem with significant computational efficiency compared to classical methods that scales with the size of the network.
@@
8.1 Theoretical and Applied Implications to Biological and Artificial Neural Networks
The importance of this work lies in both its theoretical and practical implications. For neuroscience, it provides a novel approach to model and understand the temporal and structural properties of neural networks. More broadly, it has potential applications for analyzing how brain dynamics relate to disorders associated with network disconnections, such as epilepsy, neurodevelopmental disorders such as Autism Spectrum Disorder, and neural cognitive and degenerative disorders. In machine learning and artificial intelligence, it is potentially relevant to constructing and optimizing artificial neural network models. For quantum computing, this work serves as a proof-of-concept for how existing quantum algorithms can be adapted and extended to solve applied, non-trivial interdisciplinary problems, motivating the continued exploration and expansion of quantum algorithms and computing.
Of particular relevance are the comments we expanded on in section 7.2.1 above. An extension of the ideas we explore here could lead to a novel analog-digital neural dynamic model that more closely resembles how we know biological neurons compartmentalize processes and compute. As we briefly introduced, biological neurons are not just purely digital computational units but rather highly complex analog-digital computing elements individually that collectively contribute to network dynamics. An approach that leverages the inherent properties of the Bloch sphere or an evolving Hamiltonian before collapsing to a binary output could be a unique way to think about an analog-digital hybrid neural dynamic model. Such a model could conceivably have applicability to understanding how the biological brain uniquely represents and processes information while also forming the basis of new forms of quantum machine learning and artificial neural networks. In such a model, the evolution of the network may not be observable.
8.2 Modeling the Evolution of Network Dynamics Using Quantum Computational Methods
This work raises two broad areas in which quantum computing may contribute to the scientific study of neural networks. The first is large-scale simulations of neural dynamics across scales of spatial and temporal organization, bounded and informed by the known physiology (in the brain) and known models (for artificial intelligence and machine learning). We did not address this first question here. The second is carefully chosen and structured problems about neural dynamics that use quantum algorithms carefully and in a clever way, an example being the work in this paper. These two research directions are necessarily complementary rather than mutually exclusive or distinct.
@@ 8.2 Modeling the Evolution of Network Dynamics Using Quantum Computational Methods
As also discussed above in section 7.2.1, there may also be scenarios where, instead of a network model being simulated, a network evolves naturally according to its physical makeup, i.e., activity in the brain or internal activation patterns in artificial neural networks. In this scenario, only then are the dynamics measured or observed at some arbitrary time $T_{o}$ to probe its behavior. In this case, as we showed here, a quantum computational approach to a well-structured question about the dynamics could be the only way to answer the question. This could have real-world implications. The dynamics of the network may be inaccessible or uninterpretable before or beyond $T_{o}$. For example, in biological neural networks, experimental limitations, such as noise, finite measurement windows, or ethical considerations, may constrain observations to specific time windows.